You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130925 (29 Sep 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 Sep 24-25

Observer: E. Sani, F. Cusano
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson
Instrument: LUCI

Summary

MODS1 is currently not available due to problems with the red CCD electronics. The MODS2 head electronics box is being sent to LBTO from Columbus.
At the sunset pho sky with no wind and 41% humidity.
At the beginning of night time operations the seeing is better than 0.5 on the GCS.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Cresci_8B DSF2237b K, 210_zJHK 3h40min within the specs
Cresci_8B HIP116324 K, 210_zJHK    
Galametz_31_A1 CANDELS_UDS18 zJspec, 200_H+K 1h35min within the specs
Galametz_31_A1 HIP11809 zJspec, 200_H+K    

Details

Start using LUCI

01:34 LUCI-IMA sky flat in H, K band (some have high counts)

02:08 Slewing to standard star of Cresci_8B (problem, the telescope is not slewing and the mos arm in luci is stuked)

02:22 Slewing to standard star of Cresci_8B

02:40 executeLUCI tellB_mos_210_K1.spec (seeing 0.6")
  • error from the instrument setup: tilting the grating it does not stabilize
  • the guide star is too close to the patrol field edge. We change it and re acquire

03.03 executeLUCI tellB_mos.acq

03.12 executeLUCI tellB_mos_210_K1.spec
  • the tilting stabilization at 2.17 mum is now working
  • we acquired in open loop, and when moving from one slit to another the star falls out of the second slit. We change the guiding star.

03.27 executeLUCI tellB_mos.acq

03.33 executeLUCI tellB_mos_210_K1.spec
  • it was not a problem of open loop acquisition. the script offsets are wrong and moving to the second slit we are off of it. We move to the scientific field to do not loose more time, in the meanwhile we recalculate the offsets.

03.55 executeLUCI campoB_mos.acq luci.20130925.0101-0104

04.18 executeLUCI campoB_mos_210_K_1.spec luci.20130925.0105-0116
  • the average seeing on the GCS is 0.55
  • at the end of this integration the average seeing on the GCS is variable, but always lower than 1.0

05.25 executeLUCI campoB_mos_210_K_2.spec luci.20130925.0117-0120
  • stopped at 0120.fits relative offset wrong, we fixed the problem and execute again

05:53 executeLUCI campoB_mos_210_K_2.spec luci.20130925.0121
  • stopped at 0121.fits relative offset again wrong, we fixed the problem and execute again

06:02 executeLUCI campoB_mos_210_K_2.spec luci.20130925.0122-0133
  • offset now ok, average seeing 0.5"

07:08 executeLUCI campoB_mos_210_K_3.spec luci.20130925.0134-0145
  • the average seeing on the GCS is 0.6

08.15 executeLUCI campoB_mos_210_K_4.spec luci.20130925.0146-0153

09.01 slew to the telluric
  • executeLUCI tell3B_mos.acq luci.20130925.0154-0157

09.14 executeLUCI campoB_mos_210_K_1.spec luci.20130925.0158-065

09:19 slew to standard Galametz_31_A1
  • executeLUCI CANDELS_UDS18_A1_tell_mos.acq 0166-0169

09.31 executeLUCI CANDELS_UDS18_A1_tell_mos_zJspec.spec 0170-0172

09.33 slew to target Galametz_31_A1
  • executeLUCI CANDELS_UDS18_A1_mos.acq 0173-0176

09.55 executeLUCI CANDELS_UDS18_A1_mos_zJspec_1.spec 0177-192

11:26 executeLUCI CANDELS_UDS18_A1_mos_zJspec_2.spec 0193-0195 (stopped reached airmass 1.6)

11:46 TO closes the dome

11:57 startin calibrations

12:00 executeLUCI CANDELS_UDS18_A1_mos_200_zJspec.flat 0196-0205
  • ADU are low: 1500-4500

12:09 executeLUCI CANDELS_UDS18_A1_mos_200_zJspec.flat 0206-0215 (increase dit to 10s)

12:22 executeLUCI CANDELS_UDS18_A1_mos_200_zJspec.flat 0216-0225 (increase dit to 15s, counts ok)

12:30 executeLUCI CANDELS_UDS18_A1_mos_200_zJspec.arc 0226-0229

12:39 executeLUCI campoB_flat_K.cal 0230-0239 (low counts)

12:49 executeLUCI campoB_flat_K.cal 0240-0249 (dit to 10s, good)

12:56 executeLUCI campoB_flat_H.cal 0250-0259

13:01 executeLUCI campoB_flat_H.cal 0260-0269 (dit to 10s,good)

13.10 executeLUCI campoB_arc_K.cal 0270-0273

12:22 executeLUCI campoB_arc_H.cal 0274-0277

20:18 The SA runs the Darks
  • EndOfNight run during the moring, so LUCI is configured for darks
  • 10x 4s O2DCR SAVEMODE=Integrated luci.20130926.0001-0010
  • 6x 10s O2DCR 0011-0015
  • 3x 5s O2DCR 0016-0020
  • 2x 30s O2DCR 0021-0025
  • 1x 7s O2DCR SAVEMODE=Normal 0026-0030
  • 1x 300s MER 0031-0040

-- FeliceCusano - 25 Sep 2013
Topic revision: r11 - 29 Sep 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback